OpenStudy (anonymous):

solve the following quadratic equation by factoring: 1.) x^2+2x=o 2.) 2x^2=6x 3.) 2a^2+a =o 4.) 6m^2+9m=o 5.) b^2=1

OpenStudy (anonymous):

i'll solve one as an example since they are all the same 1. \(x^2+2x=0\) take out common factor \(x(x+2)=0\) use Null factor law \(x = 0\) or \(x+2 = 0\) \(x = -2\) or \(x = 0\)

OpenStudy (anonymous):

thanks but can you solve also the number 2.

OpenStudy (anonymous):

it's the same, just move them to one side of the equation and factorise

OpenStudy (anonymous):

ahh ok. thanks again